# FusionRF: High-Fidelity Satellite Neural Radiance Fields from Multispectral and Panchromatic Acquisitions

### v2

@@ IV-B Depth Reconstruction Comparisons
Four evaluation regions, each covering approximately $0.06$ km2, were chosen from WorldView-3 satellite acquisitions available through the 2019 IEEE GRSS Data Fusion Contest [50, 51]. Each region contains between 11 and 24 views, with between 2 and 5 images withheld for evaluation. A slight contrast adjustment was applied but no color correction or normalization was performed. An external bundle adjustment [52] method was used. All images were rendered with the multispectral modal embedding. To measure the quality of constructed site representation, depth maps are generated and compared to LIDAR scans of the scene. The average error is measured with an exclusion mask applied for areas of water.
In Figures 5 and 6, we show a comparison between the generated depth from FusionRF trained with cross-resolution kernel $F_{\Psi}$ on panchromatic and multispectral data, and that of FusionRF trained with pansharpened data from PS-Dip [70]. We show visually improved clarity in detailed regions, such as small trees and building edges. Additionally, we observe better performance in reconstructing the ground, with less noise and false objects present. Figure 4 shows a visual comparison between depth reconstruction results for a sample scene.
@@
In Table II, we show the Mean Absolute Error (MAE) calculated between the LIDAR scan and the generated depth from various ablations of FusionRF. Additionally, we include results from the most recent satellite reconstruction method from satellite images with an available codebase for reproduction and fair comparison, Sat-NeRF. We show that FusionRF is able to reduce reconstruction error by 17% compared to Sat-NeRF, the performance of FusionRF increases when sparse kernel $F_{\Psi}$ is enabled, the inclusion of both $F_{\Psi}$ and input panchromatic images is necessary to perform intrinsic pansharpening, and the increase in performance provided by image and modal embeddings. Figures 5 and 6 visually show depth comparison results for sample scenes in each dataset.
IV-C Existing Neural Image Restoration Methods
@@ V Conclusion
We present FusionRF, an advancement in satellite-based NeRF methods that removes the requirement of pansharpening in satellite image preprocessing. Our model fuses low spatial resolution multispectral and low spectral resolution panchromatic imagery directly from common observation satellites and intrinsically perform pansharpening, rendering sharp novel view images with both high spectral and high spatial resolutions.

### v3

@@ IV-B Depth Reconstruction Comparisons
Four evaluation regions, each covering approximately $0.06$ km2, were chosen from WorldView-3 satellite acquisitions available through the 2019 IEEE GRSS Data Fusion Contest (JAX) [36, 37]. Each region contains between 11 and 24 views, with between 2 and 5 images withheld for evaluation. A slight contrast adjustment was applied but no color correction or normalization was performed. An external bundle adjustment [38] method was used. All images were rendered with the multispectral modal embedding. To measure the quality of constructed site representation, depth maps are generated and compared to LIDAR scans of the scene. The average error is measured with an exclusion mask applied for areas of water.
@@
In Figures 6 and 7, we show a comparison between the generated depth from FusionRF trained with cross-resolution kernel $F_{\Psi}$ on panchromatic and multispectral data, and that of FusionRF trained with pansharpened data from PSDip [48]. We show visually improved clarity in detailed regions, such as small trees and building edges. Additionally, we observe better performance in reconstructing the ground, with less noise and false objects present. Figure 4 shows a visual comparison between depth reconstruction results for a sample scene.
Figure 8: Depth Comparison Results: The results from FusionRF+EO-NeRF optimized over multispectral and panchromatic imagery aligns closer to the LIDAR ground truth than the same model optimized over pansharpened imagery generated by PSDip, demonstrating the efficacy of our method. Areas highlighted in red are enlarged in the following row.
In Table II, we show the Mean Absolute Error (MAE) calculated between the LIDAR scan and the generated depth from various ablations of FusionRF. Additionally, we include results from a baseline method, Sat-NeRF [13]. We show that FusionRF is able to reduce reconstruction error by 17% compared to Sat-NeRF, the performance of FusionRF increases when sparse cross-resolution kernel $F_{\Psi}$ is enabled, the inclusion of both $F_{\Psi}$ and input panchromatic images is necessary to perform intrinsic pansharpening, and the increase in performance provided by image and modal embeddings. Figures 6 and 7 visually show depth comparison results for sample scenes in each dataset.
IV-C Extension to EO-NeRF
TABLE III: Comparison of EO-NeRF and FusionRF methods across JAX and IARPA datasets. Best mean result bolded.
JAX IARPA
004 068 214 260 001 002 003 Mean
EO-NeRF
1.72 1.08 1.58 1.53 2.07 2.39 1.99 1.76
FusionRF + EO-NeRF
(PSDip Pansharpened)
1.36 1.36 1.88 1.62 1.68 2.17 2.04 1.73
FusionRF + EO-NeRF
(Raw Imagery)
1.29 1.15 1.72 1.41 1.51 1.70 2.29 1.58
In order to demonstrate the flexibility of our embedding strategy and cross-resolution kernel, we adapt them to the state-of-the-art method EO-NeRF [14]. As an ablation model, we train EO-NeRF on images produced by the state-of-the-art pansharpening algorithm PSDip [48] with our novel embedding strategy. For this comparison, the original RPCs provided with the images were used for all models.
For additional scenes, three evaluation regions from the 2016 IARPA Multi-View Stereo 3D Mapping Challenge (IARPA) [49] are adopted, each covering $0.06$ km2. Each area contains between 20 and 24 distinct views with 2 to 5 views withheld for validation. A slight contrast adjustment was added but no color correction or normalization was performed. We provide a visual comparison between LIDAR scans of the scene and the depth reconstruction results of our adapted model and the ablation model in Figures 8 and 9.
Figure 9: Depth Comparison Results: Areas in red in Figure 8 are enlarged to show increased detail in the clarity of building edges and rendering of small objects.
Our comparison shows the increased clarity in depth reconstruction results when using the original multispectral and panchromatic images instead of pansharpening the images upstream of the reconstruction model. We note that the depth reconstruction is clearer in sections of fine detail such as building edges, trees, and small structures.
We calculate the mean absolute error of the adapted and ablation models against the LIDAR scans and report the results in Table III. Additionally, we include the reported EO-NeRF performance on pansharpened imagery. Our results show a 10.23% decrease in reconstruction error across all datasets, indicating a performance benefit from optimizing over raw multispectral and panchromatic imagery. This adapted model demonstrates that the proposed cross-resolution kernel and embedding strategies can be adapted to any NeRF model that operates on satellite imagery to avoid the pansharpening stage.
IV-D Timings
Using the cross-resolution kernel means that the model must also render an additional eight neighboring support rays for each pixel in the input images. Additionally, FusionRF must render twice the number of pixels as pansharpened methods, since it optimizes over multispectral and panchromatic images independently. This results in a much longer training time than other models. However, since minimal parameters are used in the cross-resolution kernel the overall parameter count is very similar to the baseline models. Additionally, the inference speed of our models are identical to the baseline models since the cross-resolution kernel is disabled during inference. The exact timings and parameter counts are shown in Table IV.
@@ IV-D Timings
TABLE IV: Model Performance: While training takes substantially longer, the inference speed and model parameter count are very near that of the baseline models.
Training Time
(per epoch)
Number of
Trainable Parameters
Inference Speed
(per image)
Sat-NeRF $\sim$1h 18min 2,523,634 $\sim$41s
FusionRF $\sim$7h 56min 2,541,117 $\sim$41s
EO-NeRF $\sim$39min 781,283 $\sim$ 35s
FusionRF + EO-NeRF $\sim$6h 18min 799,030 $\sim$35s
IV-E Existing Neural Image Restoration Methods
@@ V Conclusion
We present FusionRF, an advancement in satellite-based NeRF methods that removes the requirement of pansharpening in satellite image preprocessing. Our model fuses low spatial resolution multispectral and low spectral resolution panchromatic imagery directly from common observation satellites and intrinsically performs pansharpening, rendering sharp novel view images with both high spectral and high spatial resolutions.